OpenStudy (anonymous):

Solve the equation x-5/3 = 6 for x 15 10 25 or 20?

OpenStudy (anonymous):

Multiply the whole equation through by 3 to eliminate the unwanted denominator and solve for x\[3x-5=18\]\[3x=23\]\[x=23/3\]Check it\[\frac{ 23 }{ 3 }-\frac{ 5 }{ 3 }=\frac{ 23-5 }{ 3 }=18/3 or 6\]